# RTNinja: a generalized machine learning framework for analyzing random telegraph noise signals in nanoelectronic devices

### v2

@@ Abstract
Random telegraph noise is a prevalent variability phenomenon in nanoelectronic devices, arising from stochastic carrier exchange at defect sites and critically impacting device reliability and performance. Conventional analysis techniques often rely on restrictive assumptions or manual interventions, limiting their applicability to complex, noisy datasets. Here, we introduce RTNinja, a generalized, fully automated machine learning framework for the unsupervised analysis of random telegraph noise signals. RTNinja deconvolves complex signals to identify the number and characteristics of hidden individual sources, without requiring prior knowledge of the system. The framework comprises two modular components: LevelsExtractor, which uses Bayesian inference and model selection to denoise and discretize the signal; and SourcesMapper, which infers source configurations through probabilistic clustering and optimization. To evaluate performance, we developed a Monte Carlo simulator that generates labeled datasets spanning broad signal-to-noise ratios and source complexities; across 7000 such datasets, RTNinja consistently demonstrated high-fidelity signal reconstruction and accurate extraction of source amplitudes and activity patterns. Our results demonstrate that RTNinja offers a robust, scalable, and device-agnostic tool for random telegraph noise characterization, enabling large-scale statistical benchmarking, reliability-centric technology qualification, predictive failure modeling, and device physics exploration in next-generation nanoelectronics.
@@ I Introduction
For over six decades, the semiconductor industry has been at the forefront of technological advancements, driving exponential performance growth and transforming society at a rapid pace. This progress has been propelled by the consistent miniaturization of device dimensions, towards nanometer scale, with each new technology node [1]. However, this advancement has come with significant challenges. Each technological node has introduced new scientific obstacles, compelling researchers to develop innovative solutions to overcome these limitations. One such obstacle is random telegraph noise (RTN) which has been observed by the device research community since the 1980s [2, 3]. This noise is observed in transistors and capacitors, which are fundamental building blocks of logic and memory devices. RTN is an intrinsic noise phenomenon observed as random discrete level fluctuations in the progression of device characteristics such as drain current and gate current. As a result, RTN leads to non-ideal device behavior, degrades long-term performance, and induces failures [4, 5, 6, 7, 8, 9]. The origin of RTN is attributed to defects in the gate dielectric material, which are capable of exchanging charge carriers (electrons or holes) with the device substrate via quantum mechanical tunneling [10]. This mechanism is then observed as charge fluctuations in device characteristics. Interest in this phenomenon grew when it was investigated in flash memories, revealing current fluctuations of up to 60% and threshold voltage shifts as high as 700 mV in 90-nm technology node devices [11, 12, 13, 14]. Subsequently, over the last two decades, RTN has become a vital phenomenon to investigate for any nanoelectronic device as it serves a dual purpose: (i) an important mechanism for investigating the microscopic physics of carrier-defect interactions, while (ii) also posing as an indicator of the reliability and stochastic variability in performance. Moreover, RTN affects not only the reliability of individual devices, but also the overall reliability of circuit designs [15, 16, 17, 18]. To effectively tackle device technology issues and benchmark different technologies, the development of advanced RTN analysis frameworks is essential as this leads to robust understanding of the physical and statistical aspects.
@@
Over the years, numerous techniques have been developed to analyze RTN signals and extract information such as observed number of individual RTN sources and their corresponding activities. Typically, but not exclusively, an individual RTN source is either a single defect or a chain of defects in nanoelectronic devices. Basic analysis techniques include histograms and time lag plots, with various adaptations of time lag plots documented in the literature [19, 20, 21, 22]. More advanced approaches, such as the Canny edge detection algorithm, hidden Markov models, Kohonen network, and deep learning-assisted method, have also been developed [23, 24, 25, 26, 27, 28, 29, 30]. Some fundamental and practical limitations encountered across these techniques, though not necessarily present in every single one, include:
@@
(i) applicable only for simple signals with very high signal-to-noise ratios (SNR) with reasonable sampling of RTN fluctuations
(ii) extracted information limited to observed levels and not individual sources
@@
(iii) applicable to specific cases of RTN signals, such as activity of single defects with exponentially distributed activity
@@
(iv) analysis limited to signals involving limited sources as complexity of more than two sources is not feasible
@@
(v) manual interaction is required, meaning the number of sources and their activity distribution parameters are user-defined inputs. This is extremely challenging for signals with more than two sources or for signals with low SNR or for analyzing datasets at large scale
@@
(vi) artefacts are introduced since some techniques must account for all possible levels originating from observable sources, even if some levels are not observed during the measurement window
@@
With recent rapid advancements in data analytics and machine learning, which are beneficial across various scientific and technological domains, the analysis of RTN signals can be classified as an unsupervised learning problem of descriptive nature. Descriptive type of ML algorithms are designed to conduct exploratory data analysis and extract meaningful information (in-depth characteristics). This contrasts with commonly used predictive ML algorithms, which make predictions about future outputs based on unseen inputs after training on a set of known inputs and outputs. In the context of RTN signals, the observable data typically represents the progression of a complex signal as a function of time, which is a convolution of unknown sources, their activities, and intrinsic experimental noise. In this work, we propose a machine learning (ML) framework—RTNinja that can conduct automated data-driven analysis of RTN signals. As shown in Figure LABEL:fig_focus, an observable RTN signal is provided as the input information and RTNinja extracts the hidden information comprising of individual sources along with their corresponding activity. Compared to existing RTN analysis techniques, RTNinja has the following major advantages:
(i) it can deconvolve complex signals constituting up to seven or more individual sources (subject to SNR resolution)
(ii) the process is completely automated where decision-making is primarily driven by data, without the need to predefine the number of sources or their activity distribution parameters
@@
RTNinja is a generalized framework as it can analyze RTN signals of any nanoelectronic device since no tailor-made assumptions are taken about the underlying sources and their activity distribution parameters. In the following sections, different modules of the RTNinja framework are explained in detail by means of an example dataset. Furthermore, robustness of the framework is discussed by applying it on large number of synthetic data (and therefore labeled) with varying SNR levels and data complexity. Finally, given that no ML framework can be invincible, a comprehensive outline of the limitations and scope of application is provided.
@@ II Monte Carlo-based RTN Simulator
RTN signals analysis exemplifies an unsupervised learning problem, where there is no availability of ground truth labels for the input data. In such scenarios, it is challenging to objectively evaluate the performance of an ML framework in identifying underlying patterns or structures in the data. Researchers frequently rely on heuristic methods, such as clustering quality metrics or visual inspection, to assess the performance of ML frameworks [31]. These methods are less reliable and can lead to subjective evaluations, unlike in supervised learning problems where the ground truth labels are available, which enables the direct comparison between ground truth and predicted output, facilitating a more straightforward and objective performance evaluation. In the context of RTN signals, it is possible to convert this unsupervised learning problem into a supervised learning scenario. It is a common practice in device research community to generate physics-based or statistical simulations of RTN signals, especially for individual defects studies, with predefined characteristics [10, 22, 26, 28, 30, 32]. In this work, we adopted a practice of building a Monte Carlo (MC) simulator that generates synthetic RTN signals, with all underlying attributes sampled from specific statistical distributions, as illustrated in Figure LABEL:fig_mcsim. The choice of these distributions aims at mimicking common research studies in the device research community, particularly those focused on understanding the impact of single defects on transistor characteristics. However, the RTNinja framework is designed to be agnostic to the underlying distributions. Figure LABEL:fig_mcsimexample presents an example dataset generated by the MC-based RTN simulator, which is used to explain the RTNinja ML framework.
@@
(a) (a) Schematic illustration of RTN signal analysis, a descriptive type and unsupervised learning problem, using the RTNinja machine learning framework. Input to the framework is a convoluted RTN signal resulting from the activity of a hidden number of individual sources, along with intrinsic noise fluctuations. Without any predefined estimates of the underlying information, RTNinja analyzes the signal and outputs the total number of individual sources and their corresponding activity. Ideally, the output is an exact reproduction of the input sources ($N=N^{\prime}$) with their corresponding activity. (b) Monte Carlo-based RTN Simulator. Number of sources (NN) is sampled from a Poisson distribution, where each source has three attributes: amplitude ($\Delta^{n}$), on-state activity, and off-state activity. Both state activities follow exponential distributions with the primary distribution parameters, on-state mean value ($\langle\tau_{on}^{n}\rangle$) and off-state mean value ($\langle\tau_{off}^{n}\rangle$), are sampled from a log-normal distribution. Extrinsic noise is added to the RTN signal in proportion to the overall signal magnitude. (c) The example dataset is simulated for a total time of $10^{3}$ units (e.g. seconds) at a sampling rate of 50 samples per unit. This particular signal constitutes four sources, with amplitudes within the range ($4\times 10^{-1}$33) and mean activities within the range ($3\times 10^{-1}$$5\times 10^{2}$). (d) Workflow of RTNinja machine learning framework. Observable information i.e., an experimental RTN signal, is provided as input data to the framework. RTNinja analyzes this information and outputs the hidden individual RTN sources and their corresponding progression characteristics. The framework consists of two modules: LevelsExtractor and SourcesMapper. LevelsExtractor identifies discrete constant levels and builds an optimal model of the input data. SourcesMapper finds the most likely solution i.e., number of sources along with their amplitudes and activities, that explains the optimal model produced by LevelsExtractor. Each module consists of three sub-modules, with the objectives and underlying approaches of each sub-module clearly outlined.
@@ III.1 Overview
The workflow representation of RTNinja ML framework is illustrated in Figure LABEL:fig_mlflow. RTNinja constitutes of two modules, namely LevelsExtractor and SourcesMapper.
The LevelsExtractor module is designed to optimally model observable input data by identifying transitions between discrete constant levels. This process involves three key sub-modules. Initially, an algorithm is designed based on Bayesian inference to determine a set of discrete levels where each level is characterized by a Gaussian distribution defined by its mean and standard deviation. Following this, the input data is reconstructed to reflect transitions between these levels, effectively filtering outliers or experimental variations through a de-noising condition. The combined application of these two sub-modules results in a characteristic feature model in which the input data is represented by an integer number of levels and the transitions between them. This model is refined based on a standard deviation value, which is estimated a priori from the data and used as an input for the Bayesian algorithm. Multiple feature models are generated by varying the standard deviation value, and the optimal model is selected using the Bayesian information criterion (BIC), ensuring the most accurate representation of the signal evolution.
The SourcesMapper module aims at identifying the optimal set of $N$ sources that can physically explain the observed feature model. To achieve this, the module addresses the challenge of associating each observed level with a specific state configuration of the individual sources, where each individual source is treated as a two-state entity comprising of an on-state and an off-state. A source’s amplitude is defined by the difference between the mean values of two levels, while a state configuration refers to a superposition of the states of $N$ sources. The process is carried out by three sub-modules. Firstly, the available information from LevelsExtractor, including the list of possible differences between all levels and the transitions between levels based on the quantized data, is utilized. Markov modeling is employed to extract the underlying transition probability matrix, and affinity propagation (AP) clustering is used to generate multiple $N$-source sets, referred to as candidate solutions. Secondly, for each candidate solution, the optimal baseline is estimated using a maximum likelihood method to align the state configurations with the Bayesian levels. Lastly, a custom-designed cost function is developed to identify the optimal solution from all candidate solutions. The modules and their underlying sub-modules are explained in detail in the following sections using an example dataset.
@@ III.2 LevelsExtractor
The LevelsExtractor module is designed to effectively capture and model the transitions within observable input data by identifying discrete constant levels. Bayesian algorithm and De-noiser sub-modules are applied together to construct a characteristic feature model of the input signal. By systematically varying the standard deviation parameter, estimated a priori from the data, multiple feature models are generated. To ensure the most accurate representation of the signal evolution, the optimal model is selected based on the Bayesian Information Criterion (BIC), which balances model complexity with data fidelity.
@@
Figures LABEL:fig_bayesian and LABEL:fig_noisefilter illustrate the flowcharts for the Bayesian algorithm and the De-noiser, respectively. The Bayesian algorithm initiates with the formulation of a hypothesis, starting with a single constant level, which is modeled as a Gaussian distribution characterized by two parameters: mean ($\mu$) and standard deviation ($\sigma$). These parameters are initialized using the first data point and a priori estimated standard deviation value, with the prior probability set at 0.5. The next observed data point is then used to update the hypothesis through Bayes’ theorem, which in turn updates the prior probability via the likelihood function. This function evaluates how well the observed data fits the present hypothesis, and by combining it with the prior probability, the posterior probability is obtained. Depending on whether the posterior probability exceeds or falls below a predefined threshold $P_{threshold}$ (i.e., a hyperparameter), the belief is updated either by adding a new level (if below threshold) or by refining the existing hypothesis with updating the model parameters based on new evidence. This iterative approach leverages the posterior probability from each step as the prior for the next, continually improving parameter estimates and enhancing the understanding of $\mu$ and $\sigma$. Figure LABEL:fig_levels demonstrates the application of the Bayesian algorithm on the example dataset, revealing that 11 levels are required to explain the full dataset, with four levels being predominant while the remaining levels have fewer data points.
@@
(b) Working details and results of the LevelsExtractor module. (a) Flowchart of the Bayesian algorithm, which identifies the most suitable hypothesis (number of levels and their model parameters) based on observed data using Bayes’ theorem. The hypothesis is updated if the posterior probability falls below a user-defined threshold ($10^{-15}$ in this case). The conditional probability, $P(d_{t}\mid H)$, is calculated using a custom CDF-based proximity estimator, representing the likelihood of a data point belonging to its nearest level ($P=1$ if $d_{t}=\mu_{i}$). (b) Flowchart of the De-noiser, which assigns data points to their nearest levels, applying a custom continuity condition (hyperparameter set to $c$ = 3 data points) to ensure statistical significance for transitions identified as RTN signatures. (c) Bayesian algorithm extracts 11 levels, with the total weighted distribution of these levels (red) overlapping the dataset histogram (blue), and $\mu$ values of levels represented by dashed lines. (d) Input data is successfully reconstructed, with $\mu$ values of the 11 levels represented by dashed lines. (e) Standard deviation is estimated a priori using a median moving range method ($\sigma_{MMR}$). While a minority of the input data contains the RTN transition information, remaining portion corresponds to experimental noise at a constant level, leading to $\sigma_{MMR}$ estimation. Multiple feature models are built by varying the initial $\sigma$ value input to the Bayesian algorithm. BIC values are computed for all models, where the optimal $\sigma$ value and the corresponding feature model are determined by the lowest BIC value. This does not necessarily coincide with the $\sigma_{MMR}$ since fast and small-amplitude RTN sources might remain undetectable.
@@
The levels extracted from the Bayesian algorithm are then used to reconstruct the input signal, with a hyperparameter acting as a constraint to filter out noise. The De-noiser flowchart begins by revisiting the first two data points and assigning them to their closest Bayesian level. A ’continuity’ condition is imposed, whereby a transition between levels is considered a ’true’ RTN only if a predefined number of data points $c$ (i.e., a hyperparameter) belong to the same level after the transition. This condition serves as a qualitative measure of stochasticity, as the likelihood of a transition being an outlier is low if three data points ($c$ = 3) are observed close to the same level post-transition. It is important to note that any hidden source with faster RTN activity will also be considered noise, and such data points will be forcefully assigned to the present level as if no transition occurred. The time resolution of activity that can be extracted is limited by this condition. For example, if the measurement sampling time is 20 ms, the fastest activity that can be captured is 60 ms, given the continuity constraint is set to three data points ($c$ = 3). This iterative process is used to reconstruct the full RTN signal (Figure LABEL:fig_levelsrecon).
@@
$\text{where,}\quad\log(\hat{L})=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\ln(\sigma^{2})-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{2\sigma^{2}}$

### v3

@@ Abstract
Random telegraph noise is a prevalent variability phenomenon in nanoelectronic devices, arising from stochastic carrier exchange at defect sites and critically impacting device reliability and performance. Conventional analysis techniques often rely on restrictive assumptions or manual interventions, limiting their applicability to complex, noisy datasets. Here, we introduce RTNinja, a generalized, fully automated machine learning framework for the unsupervised analysis of random telegraph noise signals. RTNinja deconvolves complex signals to identify the number and characteristics of hidden individual sources without requiring prior knowledge of the system. The framework comprises two modular components: LevelsExtractor, which uses Bayesian inference and model selection to denoise and discretize the signal, and SourcesMapper, which infers source configurations through probabilistic clustering and optimization. To evaluate performance, we developed a Monte Carlo simulator that generates labeled datasets spanning broad signal-to-noise ratios and source complexities; across 7000 such datasets, RTNinja consistently demonstrated high-fidelity signal reconstruction and accurate extraction of source amplitudes and activity patterns. Our results demonstrate that RTNinja offers a robust, scalable, and device-agnostic tool for random telegraph noise characterization, enabling large-scale statistical benchmarking, reliability-centric technology qualification, predictive failure modeling, and device physics exploration in next-generation nanoelectronics.
@@ I Introduction
For over six decades, the semiconductor industry has been at the forefront of technological advancements, driving exponential performance growth and transforming society at a rapid pace. This progress has been propelled by the consistent miniaturization of device dimensions, toward the nanometer scale, with each new technology node [cmosscale]. However, this advancement has come with significant challenges. Each technological node has introduced new scientific obstacles, compelling researchers to develop innovative solutions to overcome these limitations. One such obstacle is random telegraph noise (RTN) which has been observed by the device research community since the 1980s [ralls1984, uren1985]. This noise is observed in transistors and capacitors, which are fundamental building blocks of logic and memory devices. RTN is an intrinsic noise phenomenon observed as random discrete level fluctuations in the progression of device characteristics, such as drain current and gate current. As a result, RTN leads to non-ideal device behavior, degrades long-term performance, and induces failures [yaney1987meta, mueller1996conductance, mueller1997statistics, fukuda2007random, qiu2015impact, goda2015time]. The origin of RTN is attributed to defects in the gate dielectric material that are capable of exchanging charge carriers (electrons or holes) with the device substrate via quantum mechanical tunneling [grasser2020noise]. This mechanism is then observed as charge fluctuations in device characteristics. Interest in this phenomenon grew when it was investigated in flash memories, revealing current fluctuations of up to 60% and threshold-voltage shifts as high as 700 mV in 90-nm technology node devices [fantini2007, tega2006anomalously, kurata2007random, spinelli2021random]. Subsequently, over the last two decades, RTN has become a vital phenomenon to investigate for any nanoelectronic device as it serves a dual purpose: (i) an important mechanism for investigating the microscopic physics of carrier-defect interactions, while (ii) also posing as an indicator of the reliability and stochastic variability in performance. Moreover, RTN affects not only the reliability of individual devices, but also the overall reliability of circuit designs [kaczer2016defect, simoen2011random, matsumo2012, luo2015]. To effectively tackle device technology issues and benchmark different technologies, the development of advanced RTN analysis frameworks is essential, as this leads to a robust understanding of the physical and statistical aspects.
@@
Over the years, numerous techniques have been developed to analyze RTN signals and extract information such as the observed number of individual RTN sources and their corresponding activities. Typically, but not exclusively, an individual RTN source is either a single defect or a chain of defects in nanoelectronic devices. Basic analysis techniques include histograms and time-lag plots, with various adaptations of time-lag plots documented in the literature [yuzhelevski2000random, nagumo2009new, nagumo2010statistical, martin2014new]. More advanced approaches, such as the Canny edge detection algorithm, hidden Markov models, Kohonen networks, and deep-learning-assisted methods, have also been developed [stampfer2018characterization, stampfer2020semi, miki2012statistical, puglisi2014factorial, grill2019electrostatic, xiao2024new, gonzalez2020neural, xu2024deep]. Some fundamental and practical limitations encountered across these techniques, though not necessarily present in every single one, include the following:
@@
(i) Applicable only for simple signals with very high signal-to-noise ratios (SNR) with reasonable sampling of RTN fluctuations.
(ii) Extracted information limited to observed levels and not individual sources.
@@
(iii) Applicable to specific cases of RTN signals, such as the activity of single defects with exponentially distributed activity.
@@
(iv) Analysis limited to signals involving limited sources, as complexity of more than two sources is often not feasible.
@@
(v) Manual interaction is required, meaning the number of sources and their activity distribution parameters are user-defined inputs. This is extremely challenging for signals with more than two sources, for signals with low SNR, or for analyzing datasets at large scale.
@@
(vi) Artifacts are introduced since some techniques must account for all possible levels originating from observable sources, even if some levels are not observed during the measurement window.
@@
With recent rapid advancements in data analytics and machine learning, which are beneficial across various scientific and technological domains, the analysis of RTN signals can be classified as an unsupervised learning problem of a descriptive nature. Descriptive types of ML algorithms are designed to conduct exploratory data analysis and extract meaningful information (in-depth characteristics). This contrasts with commonly used predictive ML algorithms, which make predictions about future outputs based on unseen inputs after training on a set of known inputs and outputs. In the context of RTN signals, the observable data typically represent the progression of a complex signal as a function of time, which is a convolution of unknown sources, their activities, and intrinsic experimental noise. In this work, we propose a machine learning (ML) framework—RTNinja—that can conduct automated data-driven analysis of RTN signals. As shown in Figure LABEL:fig_focus, an observable RTN signal is provided as the input information and RTNinja extracts the hidden information comprising individual sources along with their corresponding activity. Compared to existing RTN analysis techniques, RTNinja has the following major advantages:
(i) It can deconvolve complex signals constituting up to seven or more individual sources (subject to SNR resolution).
(ii) The process is completely automated, where decision-making is primarily driven by data, without the need to predefine the number of sources or their activity distribution parameters.
@@
RTNinja is a generalized framework as it can analyze RTN signals of any nanoelectronic device since no tailor-made assumptions are taken about the underlying sources and their activity distribution parameters. In the following sections, different modules of the RTNinja framework are explained in detail by means of an example dataset. Furthermore, the robustness of the framework is discussed by applying it to a large number of synthetic (and therefore labeled) data with varying SNR levels and data complexity. Finally, given that no ML framework can be invincible, a comprehensive outline of the limitations and scope of application is provided.
@@ II Monte Carlo-based RTN Simulator
RTN signal analysis exemplifies an unsupervised learning problem, where there is no availability of ground-truth labels for the input data. In such scenarios, it is challenging to objectively evaluate the performance of an ML framework in identifying underlying patterns or structures in the data. Researchers frequently rely on heuristic methods, such as clustering quality metrics or visual inspection, to assess the performance of ML frameworks [fergus2022performance]. These methods are less reliable and can lead to subjective evaluations, unlike in supervised learning problems where the ground-truth labels are available, which enables the direct comparison between ground truth and predicted output, facilitating a more straightforward and objective performance evaluation. In the context of RTN signals, it is possible to convert this unsupervised learning problem into a supervised learning scenario. It is a common practice in the device research community to generate physics-based or statistical simulations of RTN signals, especially for individual defect studies, with predefined characteristics [grasser2020noise, martin2014new, puglisi2014factorial, xiao2024new, xu2024deep, vecchi2023unified]. In this work, we adopted the practice of building a Monte Carlo (MC) simulator that generates synthetic RTN signals, with all underlying attributes sampled from specific statistical distributions, as illustrated in Figure LABEL:fig_mcsim. The choice of these distributions aims at mimicking common research studies in the device research community, particularly those focused on understanding the impact of single defects on transistor characteristics. However, the RTNinja framework is designed to be agnostic to the underlying distributions. Figure LABEL:fig_mcsimexample presents an example dataset generated by the MC-based RTN simulator, which is used to explain the RTNinja ML framework.
@@
(a) (a) Schematic illustration of RTN signal analysis, a descriptive type of unsupervised learning problem, using the RTNinja machine learning framework. Input to the framework is a convoluted RTN signal resulting from the activity of a hidden number of individual sources, along with intrinsic noise fluctuations. Without any predefined estimates of the underlying information, RTNinja analyzes the signal and outputs the total number of individual sources and their corresponding activity. Ideally, the output is an exact reproduction of the input sources ($N=N^{\prime}$) with their corresponding activity. (b) Monte Carlo-based RTN simulator. The number of sources (NN) is sampled from a Poisson distribution, where each source has three attributes: amplitude ($\Delta^{n}$), on-state activity, and off-state activity. Both state activities follow exponential distributions with the primary distribution parameters, on-state mean value ($\langle\tau_{on}^{n}\rangle$) and off-state mean value ($\langle\tau_{off}^{n}\rangle$), are sampled from a log-normal distribution. Extrinsic noise is added to the RTN signal in proportion to the overall signal magnitude. (c) The example dataset is simulated for a total time of $10^{3}$ units (e.g., seconds) at a sampling rate of 50 samples per unit. This particular signal constitutes four sources, with amplitudes within the range ($4\times 10^{-1}$33) and mean activities within the range ($3\times 10^{-1}$$5\times 10^{2}$). (d) Workflow of the RTNinja machine learning framework. Observable information, that is, an experimental RTN signal, is provided as input data to the framework. RTNinja analyzes this information and outputs the hidden individual RTN sources and their corresponding progression characteristics. The framework consists of two modules: LevelsExtractor and SourcesMapper. LevelsExtractor identifies discrete constant levels and builds an optimal model of the input data. SourcesMapper finds the most likely solution, that is, the number of sources along with their amplitudes and activities, that explains the optimal model produced by LevelsExtractor. Each module consists of three sub-modules, with the objectives and underlying approaches of each sub-module clearly outlined.
@@ III.1 Overview
The workflow representation of the RTNinja ML framework is illustrated in Figure LABEL:fig_mlflow. RTNinja consists of two modules, namely, LevelsExtractor and SourcesMapper.
The LevelsExtractor module is designed to optimally model observable input data by identifying transitions between discrete constant levels. This process involves three key sub-modules. Initially, an algorithm based on Bayesian inference is designed to determine a set of discrete levels where each level is characterized by a Gaussian distribution defined by its mean and standard deviation. Following this, the input data are reconstructed to reflect transitions between these levels, effectively filtering outliers or experimental variations through a de-noising condition. The combined application of these two sub-modules results in a characteristic feature model in which the input data are represented by an integer number of levels and the transitions between them. This model is refined based on a standard deviation value, which is estimated a priori from the data and used as an input for the Bayesian algorithm. Multiple feature models are generated by varying the standard deviation value, and the optimal model is selected using the Bayesian information criterion (BIC), ensuring the most accurate representation of the signal evolution.
The SourcesMapper module aims at identifying the optimal set of $N$ sources that can physically explain the observed feature model. To achieve this, the module addresses the challenge of associating each observed level with a specific state configuration of the individual sources, where each individual source is treated as a two-state entity comprising an on-state and an off-state. A source’s amplitude is defined by the difference between the mean values of two levels, while a state configuration refers to a superposition of the states of $N$ sources. The process is carried out by three sub-modules. First, the available information from LevelsExtractor, including the list of possible differences between all levels and the transitions between levels based on the quantized data, is utilized. Markov modeling is employed to extract the underlying transition probability matrix, and affinity propagation (AP) clustering is used to generate multiple $N$-source sets, referred to as candidate solutions. Second, for each candidate solution, the optimal baseline is estimated using a maximum likelihood method to align the state configurations with the Bayesian levels. Finally, a custom-designed cost function is developed to identify the optimal solution from all candidate solutions. The modules and their underlying sub-modules are explained in detail in the following sections using an example dataset.
@@ III.2 LevelsExtractor
The LevelsExtractor module is designed to effectively capture and model the transitions within observable input data by identifying discrete constant levels. The Bayesian algorithm and De-noiser sub-modules are applied together to construct a characteristic feature model of the input signal. By systematically varying the standard deviation parameter, estimated a priori from the data, multiple feature models are generated. To ensure the most accurate representation of the signal evolution, the optimal model is selected based on the Bayesian Information Criterion (BIC), which balances model complexity with data fidelity.
@@
Figures LABEL:fig_bayesian and LABEL:fig_noisefilter illustrate the flowcharts for the Bayesian algorithm and the De-noiser, respectively. The Bayesian algorithm initiates with the formulation of a hypothesis, starting with a single constant level, which is modeled as a Gaussian distribution characterized by two parameters: mean ($\mu$) and standard deviation ($\sigma$). These parameters are initialized using the first data point and the a priori estimated standard deviation value, with the prior probability set at 0.5. The next observed data point is then used to update the hypothesis through Bayes’ theorem, which in turn updates the prior probability via the likelihood function. This function evaluates how well the observed data fits the present hypothesis, and by combining it with the prior probability, the posterior probability is obtained. Depending on whether the posterior probability exceeds or falls below a predefined threshold $P_{threshold}$ (i.e., a hyperparameter), the belief is updated either by adding a new level (if below the threshold) or by refining the existing hypothesis by updating the model parameters based on new evidence. This iterative approach leverages the posterior probability from each step as the prior for the next, continually improving parameter estimates and enhancing the understanding of $\mu$ and $\sigma$. Figure LABEL:fig_levels demonstrates the application of the Bayesian algorithm on the example dataset, revealing that 11 levels are required to explain the full dataset, with four levels being predominant while the remaining levels have fewer data points.
@@
(b) Working details and results of the LevelsExtractor module. (a) Flowchart of the Bayesian algorithm, which identifies the most suitable hypothesis (number of levels and their model parameters) based on observed data using Bayes’ theorem. The hypothesis is updated if the posterior probability falls below a user-defined threshold ($10^{-15}$ in this case). The conditional probability, $P(d_{t}\mid H)$, is calculated using a custom CDF-based proximity estimator, representing the likelihood of a data point belonging to its nearest level ($P=1$ if $d_{t}=\mu_{i}$). (b) Flowchart of the De-noiser, which assigns data points to their nearest levels, applying a custom continuity condition (hyperparameter set to $c$ = 3 data points) to ensure statistical significance for transitions identified as RTN signatures. (c) The Bayesian algorithm extracts 11 levels, with the total weighted distribution of these levels (red) overlapping the dataset histogram (blue), and $\mu$ values of levels represented by dashed lines. (d) Input data are successfully reconstructed, with $\mu$ values of the 11 levels represented by dashed lines. (e) Standard deviation is estimated a priori using a median moving range method ($\sigma_{MMR}$). While a minority of the input data contains the RTN transition information, the remaining portion corresponds to experimental noise at a constant level, leading to $\sigma_{MMR}$ estimation. Multiple feature models are built by varying the initial $\sigma$ value input to the Bayesian algorithm. BIC values are computed for all models, where the optimal $\sigma$ value and the corresponding feature model are determined by the lowest BIC value. This does not necessarily coincide with $\sigma_{MMR}$ since fast and small-amplitude RTN sources might remain undetectable.
@@
The levels extracted from the Bayesian algorithm are then used to reconstruct the input signal, with a hyperparameter acting as a constraint to filter out noise. The De-noiser flowchart begins by revisiting the first two data points and assigning them to their closest Bayesian level. A ’continuity’ condition is imposed, whereby a transition between levels is considered a ’true’ RTN only if a predefined number of data points $c$ (i.e., a hyperparameter) belong to the same level after the transition. This condition serves as a qualitative measure of stochasticity, as the likelihood of a transition being an outlier is low if three data points ($c$ = 3) are observed close to the same level post-transition. It is important to note that any hidden source with faster RTN activity will also be considered noise, and such data points will be forcefully assigned to the present level as if no transition occurred. The time resolution of activity that can be extracted is limited by this condition. For example, if the measurement sampling time is 20 ms, the fastest activity that can be captured is 60 ms, given the continuity constraint is set to three data points ($c$ = 3). This iterative process is used to reconstruct the full RTN signal [Figure LABEL:fig_levelsrecon].
@@
$\text{where,}\quad\log(\hat{L})=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\ln(\sigma^{2})-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{2\sigma^{2}},$